OpenStudy (anonymous):

You throw a football straight up into the air with a speed of 12.2 m/s. How high does the ball go?

OpenStudy (jdoe0001):

well, if the speed is kept constant, 12.2m/s then the football will just go off the atmosphere and off to space

OpenStudy (anonymous):

0.62 m B. 3.2 m C. 9.1 m D. 7.6 m Those are the options

OpenStudy (jdoe0001):

\(\bf \text{initial velocity}\\ h = -4.9t^2+v_ot+h_o \qquad \text{in meters}\\ \quad \\ h_o=\textit{initial height, in this case, 0}\\ v_o=\textit{initial velocity or speed, in this case 12.2}\\ \quad \\ h = -4.9t^2+v_ot+h_o \implies h = -4.9t^2+12.2t+0\) notice is a parabola, with a negative leading coefficient, thus is opening downwards, highest point is its vertex you can find the vertex of a \(\bf y=ax^2+bx+c\) equation at the coordinates of \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (jdoe0001):

the height will be the y-coordinates of course

OpenStudy (anonymous):

Which one is the correct answer ?